OpenStudy (anonymous):

pH = −log of the concentration of protons (H+ in moles per liter or molar.) A solution has a H+ concentration of 3.16 × 10−9 M. What is the pH of the solution? I feel like I must be typing this in my calculator wrong.

OpenStudy (lfreeman1285):

Try putting it in as the log(3.16 x 10^-9), and then making that number negative.

OpenStudy (irishboy123):

it's \(log_{10}\), if that helps

OpenStudy (butterflydreamer):

your equation to find pH is given by: \[pH = -\log _{10} \left[ H^+ \right]\] So we're given \[\left[ H^+ \right] = 3.16 \times 10^{-9}\] Just plug into the equation and you'll get your answer :)